# OmniQuery: Contextually Augmenting Captured Multimodal Memory to Enable Personal Question Answering

### v1

@@ 3.3. Data Summary and analysis
These queries aim to get direct answers that can be retrieved by searching for memories via description (e.g., “skateboarding in a tie-dye shirt”) or rely on information explicitly contained within a single captured instance (e.g., “What is my driver’s license number?”). Typically, this type of query does not require additional context not contained in a single captured memory.
@@
These queries are more complex, combining both direct content queries and contextual filters. For example, a participant asked “Which meat did I order the last time I came to this Japanese BBQ restaurant?” Answering such a query typically requires a multi-hop process: (1) filter all captured memories under the specific context (e.g., dining in this Japanese restaurant); (2) analyze the filtered data to generate the final result.
@@ 3.4. Analysis
To identify this essential contextual information, two researchers on the team independently analyzed the logged queries, and coded, filtered, and categorized the types of context required to filter captured memories and better answer the queries. Their results were compared, and discrepancies in categorizations, hierarchy, naming, and granularity were discussed and resolved.
@@ 4. TAXONOMY OF CONTEXTUAL INFORMATION
In this section, we present the taxonomy built from analyzing user queries. We identified three key types of contextual information that can be integrated with captured memories. These include (1) atomic context, (2) composite context and (3) semantic knowledge.
Figure 2. Number of appearance of each types of context (atomic and composite) in the logged queries. Note that a query may contain multiple types of categories, such as “What boba tea did I drink last week?
@@ 4.1. Atomic Context
[…]
@@
Atomic context refers to contextual information typically obtainable from a single captured memory. This includes data directly from metadata, sensed from visual and auditory content, or inferred from the content itself. Table 1 shows the seven types of atomic contexts categorized from the queries. Among them, temporal information and geographical info can be directly obtained from the memory media’s metadata. People and visual elements typically require machine learning models or facial recognition for detection. Environment, activity, and emotion are more implicit and require reasoning based on the content (e.g., a photo of a menu may suggest the person is in a restaurant). The number of appearance of each category is shown in Figure 2.
@@ 4.2. Composite Context
Composite context is how people often remember and refer to past experiences with a single phrase, such as ”Who did I ski with in the lab retreat last year?” These contexts can range from significant events like a wedding or a conference trip to smaller incidents like hanging out with a friend or a day trip to Seattle. Specifically, composite context is defined as a combination of multiple atomic contexts. For example, the composite context “lab retreat” encompasses atomic contexts including “Feburary, 2024” (temporal), “Lake Tahoe, California” (geographical) and “hanging out with labmates” (activity).
@@ 4.3. Semantic Knowledge
In psychology theories, semantic knowledge refers to the general world knowledge that humans accumulate over time (Tulving, 2002; McRae and Jones, 2013), distinct from episodic memories, which are tied to specific experiences and events. Similarly, we can generate semantic knowledge from a user’s captured memories, providing broader insights of the user’s past experiences. For example, patterns like “Jason has a habit of going to the gym 3-4 times a week” can be inferred from multiple captured memories. Such patterns are helpful in answering queries that not necessarily require specific knowledge such as “How often do I go to the gym in April?
@@ 5.1. Step 1: Structuring Individual Captured Memories
Content of a captured memory includes the overall description of the memory as caption, visible text, and transcribed speech transcribed speech (for videos, not shown in Figure 4). Specifically, OmniQuery leverages multimodal models to process and describe the captured memory as the caption, performs optical character recognition (OCR) to recognize visible text, and uses audio-to-text models to transcribe speech.
@@
With the content processed, OmniQuery annotates each captured memory with each type of atomic context. As shown in Figure 4b, OmniQuery extracts the temporal and geographical information from the metadata and uses multimodal models to detect people and other visual elements. Then OmniQuery synthesizes the processed information and infers the environment and activities. For example, based on a photo of a sign displaying conference Wi-Fi details, OmniQuery infers that the user is likely attending a conference (activity) and is at the conference venue (environment). Note that due to the subjective matter of emotions, which often requires user input, emotion inference is excluded in current implementation.
Figure 4. An example of structuring an individual captured memory (a photo of the Wifi details of CHI 2024 conference).
@@ 5.1.1. Indexing
After each captured memory is structured, it is indexed and stored in a database. Additionally, the annotations in textual format are encoded into text embeddings to enable vector-based search during the retrieval process. In the database, each data entry represents a data entry corresponding to a captured memory, containing both the original media (e.g., photo, video) and its structured annotations in text and text embeddings.
@@ 5.2. Step 2: Identifying Composite Context
Figure 5. An example of using sliding windows to identify composite contexts and infer semantic knowledge: (a) two consecutive sliding windows; (b) composite contexts and semantic knowledge generated in each window; (c) merging results from the two windows.
As captured memories are recorded in a linear manner along a personal timeline, those interconnected through semantic contexts often cluster closely together. For example, memories related to CHI 2024 are likely to occur during the event itself. Taking advantage of this temporal proximity, OmniQuery adopts a sliding window approach to analyze potentially interconnected captured in segments, identifying composite contexts.
As shown in Figure 5a, a static window size of seven days is used in current implementation. The inference is performed via an LLM, in which the input is the structured annotations of these memories and the output is the identified composite contexts along with their start and end dates and the associated captured memories (Figure 5b). To account for cases where composite contexts are split in half, we use a step size (4 days in the current setup) smaller than the window size, allowing for overlap and comprehensive processing. For longer composite contexts (e.g., lasting more than two weeks), each segment of the context is identified separately within the sliding windows and then merged into a single composite context. Additionally, any duplicated composite contexts caused by the overlap between sliding windows are also merged to avoid redundancy (Figure 5c).
Specifically, as opposed to including detailed predefined categories (as with atomic contexts) in the prompt for LLMs, we adopt a few-shot prompting technique (Brown et al., 2020), providing examples of composite contexts summarized from the collected questions in the prompt. For detailed prompt, please refer to Appendix A.2.
@@
However, some composite contexts are explicitly mentioned in the captured memories. For example, a screenshot of a flyer may reference the upcoming ”CHI 2024” event happening next month, or a transcribed conversation might discuss a ”Hawaii trip” that took place the previous year. We leverage LLMs to differentiate between atomic contexts and composite contexts. For example, “a workout session” is identified as an activity (atomic context), whereas “CHI 2024” is recognized as a composite context which likely involves multiple interconnected atomic contexts. Such identified composite contexts are either merged with an existing composite context (e.g., if ”CHI 2024” has already been identified) or added as a new composite context if it is unique.
@@ 5.3. Step 3: Inferring Semantic Knowledge
Different from composite contexts, semantic knowledge focuses on high-level general knowledge rather than specific memory details. For example, if a chat message screenshot mentions celebrating Jason’s birthday, the inferred semantic knowledge might be ”Jason’s birthday is on [SPECIFIC DATE].” Similarly, analyzing multiple grocery shopping receipts that consistently include lactose-free milk could lead to the inference that the user is possibly lactose intolerant. Semantic knowledge is inferred in each sliding window, while also taking into account the identified composite contexts to gain higher-level understanding of the user’s past and generalized information (Figure 5b bottom). The output is a list of inferred declarative semantic knowledge independent from specific memories. The instructions provided to the model are specifically tailored to guide the inference process toward overarching patterns and trends rather than specific event details. The detailed prompt for identifying semantic knowledge can be also found in Appendix A.2. Each inferred entry of semantic knowledge is either merged with existing entries or added to the knowledge list if new.
Figure 6. The question-answering system consists of: (a) query augmentation by decomposing and inferring contextual information; (b) retrieving memories and semantic knowledge; (c) generating answers with an LLM using referenced memories. Given a query: (d) OmniQuery rewrites and decomposes it; (e) infers contextual information; (f) retrieves relevant memories; and (g) retrieves knowledge from memory and knowledge storage.
@@ 5.4. Implementation Details
To deduplicate images as people tend to capture similar content multiple times, we use CLIP (Radford et al., 2021) to encode images to embeddings and calculate the similarities between images and merge those exceed the similarity of 0.85. We use the Google Cloud Vision API [^3] https://cloud.google.com/vision/docs/ocr for OCR to detect text in images and OpenAI’s Whisper model [^4] https://github.com/openai/whisper for audio-to-text conversion. Note that Whipser is known for hallucination when there is no speech in the audio, thus we applied further data cleaning to validate the transcribed result using OpenAI’s GPT-4o-mini. For other visual processing, We use GPT-4o handles multimodal sensing, including identifying people and visual elements in images and generating scene descriptions. For video processing, as a proof-of-concept, we consider only the first 10 seconds of each video, sampling 10 frames to be analyzed by GPT-4o for content understanding. Text is encoded into embeddings using OpenAI’s text-embedding-3-small model. Currently, we utilize a custom vector database and matrix-based similarity search implemented with NumPy in Python. However, for real-world applications, more advanced vector databases (e.g., Pinecone [^5] https://www.pinecone.io/) would be necessary to handle larger volumes of personal data.
@@ 6. OMNIQUERY: QUESTION-ANSWERING SYSTEM
As shown in Figure 6, given an input query, OmniQuery first augments the query by disambiguating and decomposing it into specific contextual elements (Figure 6a). Then, it retrieves the relevant captured memories from the structured captured memories and the composite contexts, along with related knowledge from the list of semantic knowledge (Figure 6b). The retrieved memories and knowledge, along with the augmented query, are then sent to an LLM to generate a comprehensive answer (Figure 6c). The detailed implementation of each step is discussed below.
6.1. Query Augmentation
As mentioned in Section 3.3, most user queries tend to be hybrid in nature or require contextual information. This means that directly searching based solely on the content of captured memories often results in an incomplete or insufficient retrieval of relevant memories. To enhance the retrieval process, OmniQuery adopts the query refinement concept from NLP (Chan et al., 2024) to augment the queries. The query augmentation process involves
@@ 6.1. Query Augmentation
    Decomposing the query to extract necessary contextual filters, such as time, location, or events, which are grounded in the taxonomy. Note that only explicitly mentioned temporal contexts like “… last week” will be recognized temporal filters. Phrase like “… during CHI 2024” are part of a composite context, thus not counted as a temporal filter;
@@
For example, as shown in Figure 6d-g, the query “What social events did I attend during CHI 2024?” is rewritten into a declarative format of “The social events I attended during CHI 2024”. Since “CHI 2024” is explicitly mentioned and identified as a composite context, it is extracted and labeled with the appropriate composite context tag. “Social events” is also extracted and identified as an atomic context (activities). Additionally, because ”social events” might include various activities like parties, dancing, or casual conversations, and likely involve multiple people present, OmniQuery infers the relevant atomic contexts (people and activities) and annotates them in the corresponding context category.
@@ 6.2. Retrieving Relevant Augmented Memories
The decomposed augmented query is used to comprehensively retrieve from the augmented captured memories. Specifically, the augmented captured memories consist of the structured captured memories (with processed content and annotated atomic contexts), the list of identified composite contexts and the list of semantic knowledge. OmniQuery uses the decomposed components from the augmented query to perform a multi-source retrieval, pulling related memories from each of these sources. The results are then consolidated into a comprehensive list of relevant memories, which are used to generate an accurate and detailed answer for the user’s query.
Declarative query $\rightarrow$ Semantic knowledge & Processed content
The declarative query is first encoded into text embeddings and used to search both the semantic memories and processed content (caption and visible text) of the captured memories This initial search step focuses on finding knowledge and memories directly related to the input query, without incorporating additional contextual filters.
Decomposed atomic contexts $\rightarrow$ Annotated atomic contexts
Each element of the decomposed atomic contexts (both extracted or inferred) is encoded into text embeddings and searched through the corresponding categories in the structured captured memory database. For example, if the query involves activities like ”party” and ”dancing,” OmniQuery searches for captured memories annotated with similar activities. Any memories that have been annotated with related or similar activities will be retrieved, ensuring that relevant memories are included in the results. Additionally, temporal contexts apply a strict filter, excluding memories outside the specified time frame (e.g., “last month”) from the retrieval process.
Decomposed composite contexts $\rightarrow$ Identified composite contexts
Any composite context decomposed in the augmenting process is also encoded into text embeddings and searched through the list of identified composite contexts. All captured memories linked to the semantically similar composite contexts are retrieved. This ensures all memories related to the composite contexts are included. Additionally, OmniQuery leverages an LLM to assess whether a composite context includes temporal constraints. For example, “… during CHI 2024” implies a strict temporal filter, while photos related to CHI 2024” does not.

### v2

@@ 3.3. Data Summary
These queries aim to get direct answers that can be retrieved by searching for memories via description (e.g., “skateboarding in a tie-dye shirt”) or rely on information explicitly contained within a single captured instance (e.g., “What is my driver’s license number?”). This type of query does not require additional context not contained in a single captured memory.
@@
These queries are more complex, combining both direct content queries and contextual filters. For example, a participant asked “Which meat did I order the last time I came to this Japanese BBQ restaurant?” Answering such a query typically requires a multi-hop process: (1) filter all captured memories under the specific context (e.g., dining in this Japanese restaurant) and (2) analyze the filtered data to generate the final result.
@@ 3.4. Analysis
To identify this essential contextual information, two researchers on the team independently analyzed the logged queries. They coded, filtered, and categorized the types of context required to filter captured memories and better answer the queries. Their results were compared, and discrepancies in categorizations, hierarchy, naming, and granularity were discussed and resolved.
@@ 4. TAXONOMY OF CONTEXTUAL INFORMATION
In this section, we present the taxonomy built from analyzing user queries. We identified three key types of contextual information that can be integrated with captured memories: (1) atomic context, (2) composite context, and (3) semantic knowledge.
@@ 4.1. Atomic Context
row3-9=5mm,row11=7mm,column1=22mm,c,column2=68mm,c,column3=67mm,c,colspec=—Q[c]—Q[c]—Q[l]—,rowspec=—Q[m]—Q[m]—Q[m]—Q[m]—Q[m]—Q[m]—Q[m]—Q[m]—Q[m]—Q[m]—Q[m]— Category & Definition Exemplar queries     refers to contextual cues
@@
Atomic context refers to contextual information typically obtainable from a single captured memory. This includes data directly from metadata, sensed from visual and auditory content, or inferred from the content itself. Table 1 shows the seven types of atomic contexts categorized from the queries. Among them, temporal information and geographical info can be directly obtained from the memory media’s metadata. People and visual elements typically require facial recognition or other vision models for detection. Environment, activity, and emotion are more implicit and require reasoning based on the content (e.g., a photo of a menu may suggest the person is in a restaurant). The number of appearances of each category is shown in Figure 2.
Figure 2. Number of appearances of each types of context (atomic and composite) in the logged queries. Note that a query may contain multiple types of categories, such as “What boba tea did I drink last week?
@@ 4.2. Composite Context
Composite context is how people remember and refer to past experiences, such as ”Who did I ski with in the lab retreat last year?” These contexts can range from significant events like a wedding or a conference trip to smaller incidents like hanging out with a friend or a day trip to Seattle. Specifically, composite context is defined as a combination of multiple atomic contexts. For example, the composite context “lab retreat” encompasses atomic contexts including “February, 2024” (temporal), “Lake Tahoe, California” (geographical), and “hanging out with labmates” (activity).
@@ 4.3. Semantic Knowledge
In psychology theories, semantic knowledge refers to the general world knowledge that humans accumulate over time (Tulving, 2002; McRae and Jones, 2013), distinct from episodic memories that are tied to specific experiences and events. Similarly, we can generate semantic knowledge from a user’s captured memories, providing broader insights of the user’s past experiences. For example, patterns like “Jason has a habit of going to the gym 3-4 times a week” can be inferred from multiple captured memories. Such patterns are helpful in answering queries that not necessarily require specific knowledge such as “How often do I go to the gym in April?
@@ 5.1. Step 1: Structuring Individual Captured Memories
Content of a captured memory includes an overall description of the memory as caption, visible text in the image, and transcribed speech (for videos, not shown in Figure 4). Specifically, OmniQuery leverages multimodal models to process and generate image captions, performs optical character recognition (OCR) to recognize visible texts, and uses audio-to-text models to transcribe speech.
@@
With the content processed, OmniQuery annotates each captured memory with each type of atomic context. As shown in Figure 4b, OmniQuery extracts the temporal and geographical information from the metadata and uses multimodal models to detect people and other visual elements. Then OmniQuery synthesizes the processed information and infers the environment and activities. For example, based on a photo of a sign displaying conference Wi-Fi details, OmniQuery infers that the user is likely attending a conference (activity) and is at the conference venue (environment). Note that due to the subjective nature of emotions that often requires user input, emotion inference is excluded from the current implementation.
Figure 4. An example of structuring an individual captured memory (a photo of the Wi-Fi details of CHI 2024 conference).
@@ 5.1.1. Indexing
After each captured memory is structured, it is indexed and stored in a database. Additionally, the annotations in textual format are encoded into text embeddings to enable vector-based search during the retrieval process. In the database, each data entry corresponds to a captured memory with both the original media (e.g., photo, video), its structured annotations in natural language, and text embeddings.
@@ 5.2. Step 2: Identifying Composite Context
Figure 5. An example of using sliding windows to identify composite contexts and infer semantic knowledge: (a) two consecutive sliding windows; (b) composite contexts and semantic knowledge generated in each window; (c) merging results of the windows.
As captured memories are recorded in a linear manner along a personal timeline, those interconnected through semantic contexts often cluster closely together. For example, memories related to CHI 2024 are likely to occur during the event itself. Taking advantage of this temporal proximity, OmniQuery adopts a sliding window approach to analyze potentially interconnected memories scattered in segments for composite context identification.
As shown in Figure 5a, a static window size of seven days is used in our current implementation. The inference is performed via an LLM, in which the input is the structured annotations of these memories and the output is the identified composite contexts along with their start and end dates and the associated captured memories (Figure 5b). To account for cases where composite contexts are split in half, we use a step size (4 days in the current setup) smaller than the window size, allowing for overlap and comprehensive processing. For longer composite contexts (e.g., lasting more than two weeks), each segment of the context is identified separately within the sliding windows and then merged into a single composite context. Additionally, any duplicated composite contexts caused by the overlap between sliding windows are also merged to avoid redundancy (Figure 5c). Note that we determined the window size heuristically. A longer sliding window can better capture extended events or patterns, while it may underperform on shorter contexts due to redundant information. One way to optimize is to make the size dynamic, adjusting it based on the density of activities in a given period. This approach would require a fixed dataset for experiments. We further discuss this in Section 9.1.
Specifically, as opposed to including detailed predefined categories (as with atomic contexts) in the prompt for LLMs, we adopt the few-shot prompting technique (Brown et al., 2020), providing examples of composite contexts summarized from the collected questions in the prompt. For the detailed prompt, please refer to Appendix A.1.
@@
Some composite contexts are explicitly mentioned in the captured memories. For example, a screenshot of a flyer may reference the upcoming ”CHI 2024” event happening next month, or a transcribed conversation might discuss a ”Hawaii trip” that took place the previous year. We leverage LLMs’ pretrained world knowledge to differentiate between atomic contexts and composite contexts. For example, “a workout session” is identified as an activity (atomic context) because, based on world knowledge, it is more likely to refer to this activity alone. In contrast, “CHI 2024” is recognized as a composite context, as it likely involves multiple interconnected atomic contexts. Such identified composite contexts are either merged with an existing composite context (e.g., if ”CHI 2024” has already been identified) or directly added as a new composite context if it is unique.
@@ 5.3. Step 3: Inferring Semantic Knowledge
Different from composite contexts, semantic knowledge focuses on high-level general knowledge rather than specific memory details. In the scenarios of personal memory, semantic knowledge refers to personalized knowledge distilled from an individual’s past, as opposed to general knowledge (e.g., “the capital of France is Paris.”). For example, if a person’s captured memories contain photos of attending CHI 2024, the distilled semantic knowledge might be, ”The person attended the CHI conference in Honolulu in 2024.” Additionally, semantic knowledge goes beyond summarizing past events. It also encompasses inferred patterns and facts about the individual’s behavior or preferences. For example, a chat message mentioning Jason’s birthday could infer that ”Jason’s birthday is on [SPECIFIC DATE].” Similarly, analyzing multiple grocery shopping receipts that consistently include lactose-free milk could lead to the inference that the user is possibly lactose intolerant. We would like to note that, while inspired by human semantic knowledge as defined in psychology (Tulving, 2002), the semantic knowledge referenced here is slightly different. In OmniQuery’s setting, the semantic knowledge is objective and derived solely from the content of the memories. In contrast, human semantic knowledge is more implicit, containing broader associations and longer-term effects that might go beyond just the observable data.
Semantic knowledge is inferred in each sliding window, while also taking into account the identified composite contexts (in Step 2) to gain higher-level understandings of the user’s past and generalized information (Figure 5b bottom). The output is a list of inferred declarative semantic knowledge independent from specific memories. The instructions provided to the model are specifically tailored to guide the inference process toward overarching patterns and trends rather than specific event details. The detailed prompt for identifying semantic knowledge can be also found in Appendix A.2. Each inferred entry of semantic knowledge is either merged with existing entries or added to the knowledge list if new.
@@ 5.4. Implementation Details
To deduplicate images and videos as people tend to capture similar content multiple times, we use CLIP (Radford et al., 2021) to encode images (or the first frame of videos) into embeddings, calculate the similarities between images, and merge those with the similarity above 0.85. We use the Google Cloud Vision API [^4] https://cloud.google.com/vision/docs/ocr for OCR to detect text in images and OpenAI’s Whisper model [^5] https://github.com/openai/whisper for audio-to-text conversion. Note that Whipser is known for hallucination when there is no speech in the audio, thus we applied further data cleaning to validate the transcribed result using OpenAI’s GPT-4o-mini. For other visual processing, We use GPT-4o handles multimodal sensing, including identifying people and visual elements in images and generating scene descriptions. For video processing, as a proof-of-concept, we consider only the first 10 seconds of each video, sampling 10 frames to be analyzed by GPT-4o for content understanding. Text is encoded into embeddings using OpenAI’s text-embedding-3-small model. Currently, we utilize a custom vector database and matrix-based similarity search implemented with NumPy in Python. However, for real-world applications, more advanced vector databases (e.g., Pinecone [^6] https://www.pinecone.io/) would be necessary to handle larger volumes of personal data.
@@ 6. OMNIQUERY: QUESTION-ANSWERING SYSTEM
As shown in Figure 6, given an input query, OmniQuery first applies a taxonomy-based augmentation of the query by disambiguating and decomposing it into specific contextual elements (Figure 6a). Then, it retrieves the relevant captured memories from the structured captured memories and the composite contexts, along with related knowledge from the list of semantic knowledge (Figure 6b). The retrieved memories and knowledge, along with the augmented query, are then sent to an LLM to generate a comprehensive answer (Figure 6c). We discuss detailed implementations of each step below.
6.1. Taxonomy-Based Query Augmentation
Figure 6. The question-answering system consists of: (a) taxonomy-based query augmentation by decomposing and inferring contextual information; (b) retrieving memories and semantic knowledge; (c) generating answers with an LLM using referenced memories. Specifically, given an input query (d), OmniQuery first augmens it via rewriting and decomposing and inferring contextual information (e). The augmented query with different types of decomposed contextual information (f) is used to retrieve captured memories from different memory and knowledge storage (g).
As mentioned in Section 3.3, most user queries tend to be hybrid in nature or require contextual information. This means that directly searching based solely on the content of captured memories often results in an incomplete or insufficient retrieval of relevant memories. To enhance the retrieval process, OmniQuery adopts the query refinement approach (Chan et al., 2024) to augment the queries. This query augmentation process is also informed by the taxonomy of contextual information and it involves
@@ 6.1. Taxonomy-Based Query Augmentation
    Decomposing the query to extract necessary contextual filters, such as time, location, or events, which are grounded in the taxonomy. Note that only explicitly mentioned temporal contexts like “… last week” will be recognized temporal filters. Phrases like “… during CHI 2024” are part of a composite context and thus not counted as a temporal filter;
@@
For example, as shown in Figure 6d-g, the query “What social events did I attend during CHI 2024?” is rewritten into a declarative format of “The social events I attended during CHI 2024”. We leveraged an LLM to classify the mentioned contexts in the query as either atomic or composite. Detailed prompts are provided in Appendix A.3. Since “CHI 2024” is explicitly mentioned and identified as a composite context, it is extracted and labeled with the appropriate composite context tag. “Social events” is also extracted and identified as an atomic context (activities). Additionally, because “social events” might include various activities like parties, dancing, or casual conversations and involve multiple people, OmniQuery infers the relevant atomic contexts (people and activities) and annotates them in the corresponding context category.
@@ 6.2. Retrieving Relevant Augmented Memories
The decomposed augmented query is used to comprehensively retrieve relevant augmented captured memories. The augmented captured memories consist of the structured captured memories (with processed content and annotated atomic contexts), the list of identified composite contexts, and the list of semantic knowledge. OmniQuery uses the decomposed components from the augmented query to perform a multi-source retrieval, pulling related memories from each of these sources. The results are then consolidated into a comprehensive list of relevant memories, which are used to generate an accurate and detailed answer for the user’s query.
Declarative query $\rightarrow$ Semantic knowledge & processed content. The declarative query is encoded into text embeddings to search for both the semantic memories and processed content (caption and visible text) of the captured memories. This initial search step focuses on finding knowledge and memories directly related to the input query, without incorporating additional contextual filters.
Decomposed atomic contexts $\rightarrow$ Annotated atomic contexts. Each element of the decomposed atomic contexts (both extracted or inferred) is encoded into text embeddings and searched through the corresponding categories in the structured captured memory database. For example, if the query involves activities like ”party” and ”dancing,” OmniQuery searches for captured memories annotated with similar activities. Any memories that have been annotated with related or similar activities will be retrieved, ensuring that relevant memories are included in the results. Additionally, temporal contexts apply a strict filter, excluding memories outside the specified time frame (e.g., “last month”) from the retrieval process.
Decomposed composite contexts $\rightarrow$ Identified composite contexts. Any composite context decomposed in the augmenting process is also encoded into text embeddings and searched through the list of identified composite contexts. All captured memories linked to the semantically similar composite contexts are retrieved. This ensures all memories related to the composite contexts are included. Additionally, OmniQuery leverages an LLM to assess whether a composite context includes temporal constraints. For example, “… during CHI 2024” implies a strict temporal filter, while “photos related to CHI 2024” does not.